### 27 Dec 2023 v1

# Maximum Preference Optimization with Importance Sampling

Abstract
Preference learning is a key technology for aligning language models with human values. Reinforcement Learning from Human Feedback (RLHF) is a model based algorithm to optimize preference learning, which first fitting a reward model for preference score, and then optimizing generating policy with on-policy PPO algorithm to maximize the reward. The processing of RLHF is complex, time-consuming and unstable. Direct Preference Optimization (DPO) algorithm using off-policy algorithm to direct optimize generating policy and eliminating the need for reward model, which is data efficient and stable. DPO use Bradley-Terry model and log-loss which leads to over-fitting to the preference data at the expense of ignoring KL-regularization term when preference near deterministic. IPO uses a root-finding pairwise MSE loss to solve the ignoring KL-regularization problem, and learning an optimal policy. But IPO’s pairwise loss still can’t s make the KL-regularization to work. In this paper, we design a simple and intuitive off-policy preferences optimization algorithm from an importance sampling view, and add an off-policy KL-regularization term which makes KL-regularization truly effective. To simplify the learning process and save memory usage, we can generate regularization data in advance, which eliminate the needs for both reward model and reference policy in the stage of optimization.
Keywords Large Language Model  $\cdot$ Preference Learning  $\cdot$ Reinforcement Learning
1 Introduction
Large language models (LLMs) Brown et al. (2020) Chowdhery et al. (2023) Bubeck et al. (2023) Radford et al. (2019) with massive scale parameters trained on large amount of data using pretrain, supervised fine-tune(SFT)Wei et al. (2021), and instruction fine-tune (IFT) Chung et al. (2022) algorithms has lead to surprising capabilities like in few-shot in context learning. The training data comes from variety of areas and has different quality, the training algorithms (pretrain, SFT, IFT) all based on maximum likelihood estimation (MLE) which learning to match the distribution of data. LLMs trained on these data using MLE algorithm generate contents with a quality gap to human judgement or values.
Preference learning Ziegler et al. (2019) Bai et al. (2022) Christiano et al. (2017) Stiennon et al. (2020) algorithm significantly improve generate quality to align with human values. It first collects pairs of generations under the same context, and a pairwise human preference to indicate which generation is better. Then a preference learning algorithm is used to optimize generating policy to generate better candidate from the pair.
Reinforcement learning form human feedback (RLHF)Ouyang et al. (2022) uses reward-model-based reinforcement learning algorithm to learn the optimal policy. It first learns a reward model from the preference data, then using a on-policy PPO Schulman et al. (2017) algorithm to maximize the learned reward. The reward is learned to use Bradley-Terry model Bradley and Terry (1952), which assumes the preference score can be approximated from substituted with point-wise reward. This assumption may lead to an approximation error. The PPO algorithm is used on data sampled from generating policy, which may has a different support or distribution drift from preference pairwise data. So, RLHF also assumes reward model has an out-of-distribution generality. Additionally, the process to train reward model and on-policy PPO algorithm is complex, time-consuming, and unstable.
Direct preference optimization (DPO)Rafailov et al. (2023) combines off-policy algorithm and Bradley-Terry model to direct learn the generating policy from preference data. The off-policy algorithm based on KL-regularization reward maximization from off-RL community, which is data efficient, stable and eliminating the need for a reward model. DPO uses the implicit reward from off-policy algorithm to optimize the Bradley-Terry model. Because the Bradley-Terry model is used to approximate the preference score, so DPO is actually do preference score matching but not reward maximizing, which makes DPO suboptimal. Also, DPO uses Bradley-Terry model after KL-regularization, it leads to ignore the KL-regularization term and over-fitting the preference dataset.
Identity mapping preference optimization (IPO)Azar et al. (2023) also uses off-policy algorithm with KL-regularization to learn the generating policy from preference data. It learns to maximize preference probability under KL-regularization, which change DPO’s preference matching optimization to preference probability maximization, and use root finding mean square error(MSE) loss to solve the maximization problem(under KL-regularization). But it doesn’t know what does preference probability maximization actually learns theoretically, and because of the distribution or support shift between preference data distribution and policy generated data distribution, the KL-regularization may also fail.
In this paper, we design a simple and intuitive off-policy preferences optimization algorithm from an importance sampling view, and add an off-policy KL-regularization term which makes KL-regularization truly effective. Our key contribution of this paper can summary as follows:
  • formalize preference learning as a preference/reward maximization problem, and design a simple and intuitive off-policy algorithm from importance sampling view
  • figure out KL-regularization fails using pairwise preference loss, and design an off-policy sample loss to make KL-regularation truly effective
  • eliminate the needs for both reward model and reference policy to save memory usage
Figure 1: Maximum Preference Optimization (MPO) direct optimize preference maximization on preference data using off-policy algorithm, and use regularization data generated in advance to make KL-regularation truly effective, which also eliminate the needs for both reward model and reference policy in optimization stage.
2 Preliminaries
The main pipeline of preference learning usually consists of three phases: 1) pretraining and supervised fine-tuning (SFT), where SFT is not a must; 2) preference data collection; 3) reinforcement-learning optimization.
Pretraining and SFT phase
Preference learning typically started with a pretrained LLMs or LLMs fine-tuned on high quality data using maximum likelihood estimation. We define the final policy after this phase as $\pi_{\textnormal{ref}}$, and the data distribution $\pi_{\textnormal{ref}}$ to approximate as $x_{s},y_{s}\sim\mathcal{D}_{s}$, so
$\pi_{\textnormal{ref}}\approx\arg\max_{\pi}\mathbb{E}_{x_{s},y_{s}\sim\mathcal{D}_{s}}\log\pi(x_{s})\log(y_{s}|x_{s})$ (1)
Preference data collection phase
After pretraining and SFT phase, $\pi_{\textnormal{ref}}$ is prompted by context $x$, and generate two responses $y_{w},y_{l}\sim\pi_{\textnormal{ref}}(\cdot|x)$. Then $x,y_{w},y_{l}$ is labeled by humans to judge which response is preferred, and denote $y_{w}\succ y_{l}|x$ if $y_{w}$ is preferred, and $y_{l}\succ y_{w}|x$ if $y_{l}$ is preferred. We define a new symbol $I=\mathbb{I}[y_{w}\succ y_{l}|x]$, and all $<x,y_{w},y_{l},I>$ consist the preference dataset $\mathcal{D}^{p}$:
$\langle x,y_{w},y_{l},I\rangle\sim\mathcal{D}^{p}.$ (2)
We also define $\rho$ as the context distribution of $x$ and $\mu$ as the preference pair distribution given context $x$ from preference data distribution.
$x\sim\rho$ (3)
$y_{w},y_{l},I\sim\mu.$ (4)
Let $p^{*}(y_{w}\succ y_{l}|x)=\mathbb{E}_{y_{w},y_{l},I\sim\mu}[{\mathbb{I}\{I=1\}}|x]$, which denotes the preference probability of $y_{w}\succ y_{l}$ given context $x$. Then the expected preference of $y_{w}$ over $\mu$, noted $p^{*}(y\succ\mu|x)$, via the following equation:
$p^{*}(y\succ\mu|x)=\underset{y_{l}\sim\mu(\cdot|x)}{\mathbb{E}}[p^{*}(y_{w}\succ y_{l}|x)].$ (5)
For any policy $\pi$, denote the total preference of policy $\pi$ to $\mu$ as
$p^{*}(\pi\succ\mu)=\underset{x\sim\rho\atop y_{\sim}\pi(\cdot|x)}{\mathbb{E}}[p^{*}(y\succ\mu|x)].$ (6)
Reinforcement-learning optimization phase
At the final phase, a point-wise reward is learned explicitly or implicitly from the preference data, and a on-policy or off-policy learning algorithm is used to optimize the reward. In previous research, the point-wise reward is learned to use Bradley-Terry model. Given context x, define $r^{*}(y|x)$ as the reward of generating $y$. Bradley-Terry model assumes the preference probability $p^{*}(y_{w}\succ y_{l}|x)$ as:
$p^{*}(y_{w}\succ y_{l}|x)=\frac{\exp(r^{*}(y_{w}|x))}{\exp(r^{*}(y_{w}|x))+\exp(r^{*}(y_{l}|x))}=\sigma(r^{*}(y_{w}|x)-r^{*}(y_{l}|x))$ (7)
where $\sigma(\cdot)$ is the sigmoid function. The $r^{*}(y|x)$ can be estimated by a parameterized model $r_{\phi}(y|x)$ via maximum likelihood estimation.
The reward-maximization or KL-regularized reward-maximization objective is used for reinforcement learning policy optimization:
$\max_{\pi_{\theta}}\underset{{y\sim\pi_{\theta}(\cdot|x)}}{\mathbb{E}}[r_{\phi}(y|x)]-\beta\mathbb{D}_{\textnormal{KL}}[\pi_{\theta}||\pi_{\textnormal{ref}}]$ (8)
where $\mathbb{D}_{\textnormal{KL}}$ is the KL-divergence and $\beta$ is the regularization weight. This objective is optimized by on-policy REINFORCE Mnih et al. (2016) or PPO algorithm.
In this paper, we define a new pair-wise reward $r^{p}(y_{w}\succ y_{l}|x)=p^{*}(y_{w}\succ y_{l}|x)$, and design new algorithm to direct optimize the pair-wise reward maximum objective.
3 Background
3.1 Reinforcement Learning from Human Feedback (RLHF)
The RLHF uses standard two-phase reward-model based reinforcement learning to maximize the reward. It contains two step: 1) reward estimation from preference data 2) reward maximization using PPO algorithm.
Reward estimation from preference data
RLHF uses 7 to model the point-wise reward, and optimize log loss to estimate the reward. The estimated reward is defined as parameterized $r_{\phi}$, and loss function defined as:
$\mathcal{L}(r_{\phi})=-\underset{x,y_{w},y_{l},I\sim\mathcal{D}^{p}}{\mathbb{E}}[I\cdot\sigma(r_{\phi}(y_{w}|x)-r_{\phi}(y_{l}|x))+(1-I)\cdot\sigma(r_{\phi}(y_{l}|x)-r_{\phi}(y_{w}|x))],$ (9)
where $I=\mathbb{I}[y_{w}\succ y_{l}|x]$.
The loss 9 is doing maximize likelihood estimation, and the estimated reward $r_{\phi}$ is used to approximate probability $p(y_{w}\succ y_{l}|x)$ from preference data distribution $\mathcal{D}^{p}$.
Reward maximization using PPO algorithm
The second phase of RLHF is optimizing objective 8 using $r_{\phi}$ learned from 9. PPO is an on-policy algorithm which continues collect data from the current policy $\pi_{\theta}$and estimated reward model, then it uses these data to estimate the gradient of 8, and then update the current policy. Because $\pi_{\theta}$ is different from $\pi_{\textnormal{ref}}$, samples generated by $\pi_{\theta}$ has a different distribution of $D_{\textnormal{ref}}$. So, RLHF assumes $r_{\phi}$ can generalize to out-of-distribution samples generated by $\pi_{\theta}$. Following prior work, it is straightforward to show that the optimal solution of 8 takes the form:
$\pi_{\theta}\propto\pi_{\textnormal{ref}}\exp(\frac{1}{\beta}r_{\phi}(x,y)).$ (10)
RLHF needs to learn a reward model, and maximize the learned reward using on-policy PPO algorithm, this process is complex, time-consuming, and unstable.
3.2 Direct Preference Optimization (DPO)
An alternative RL algorithm to preference learning is direct preference optimization (DPO), which eliminates the training of reward model. DPO derives a reward from Eq. 10:
$r(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\textnormal{ref}}(y|x)}+\beta\log Z(x),$ (11)
where $Z(x)=\sum_{y}\pi_{\textnormal{ref}}(y|x)\exp(\frac{1}{\beta}r(x,y))$ is the partition function. Substituting the re-parameterized reward in 11 int to the Bradley-Terry model 7:
$\begin{split}p_{\theta}(y_{w}\succ y_{l}|x)&=\sigma(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)})\\ &=\frac{1}{1+\exp\big{(}\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}\big{)}}.\end{split}$ (12)
Substituting the probability 12 to loss 9, DPO formulates a maximum likelihood objective for the parameterized policy $\pi_{\theta}$ given the empirical preference dataset $\mathcal{D}^{p}$:
$\begin{split}\mathcal{L}_{\textnormal{DPO}}(\pi_{\theta};\pi_{\textnormal{ref}})=-\underset{x,y_{w},y_{l},I\sim\mathcal{D}^{p}}{\mathbb{E}}\big{[}I\cdot&\sigma(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)})\\ +(1-I)\cdot&\sigma(\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}-\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)})\big{]}.\end{split}$ (13)
In its population form, the loss takes on the form
$\begin{split}-\underset{x\sim\rho\atop y_{w},y_{l},I\sim\mu(\cdot|x)}{\mathbb{E}}\big{[}p^{*}(y_{w}\succ y_{l}|x)&\log\sigma(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)})\\ +(p^{*}(y_{l}\succ y_{w}|x))&\log\sigma(\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}-\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}))\big{]}.\end{split}$ (14)
Although this pair-wise loss eliminates the need to calculate the partition $Z(x)$, it also makes the optimal solution $\pi^{*}_{\theta}$ undefined. For example, if weight $\pi_{\theta}(y_{w}|x)$ and $\pi_{\theta}(y_{l}|x)$ with the same multiplier $M$, the logits of the sigmoid function $\sigma$ will remain the same:
$\beta\log\frac{\pi_{\theta}(y_{w}|x)*M}{\pi_{\textnormal{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)*M}{\pi_{\textnormal{ref}}(y_{l}|x)}=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}.$ (15)
This will make the final learned policy $\pi_{\theta}$ suboptimal, and also fails the KL-regularation term.
3.3 $\Psi$-PO with dientity mapping (IPO)
IPO defines a new objective called $\Psi$-preference optimization objective ($\Psi$PO):
$\max\limits_{\pi_{\theta}}\underset{\begin{subarray}{c}x\sim\rho\\ y_{w}\sim\pi_{\theta}(\cdot|x)\\ y_{l}\sim\mu(\cdot|x)\end{subarray}}{\mathbb{E}}[\Psi(p^{*}(y_{w}\succ y_{l}|x))]-\tau\mathbb{D}_{\textnormal{KL}}[\pi_{\theta}||\pi_{\textnormal{ref}}],$ (16)
where $\Psi$ is a general non-decreasing function $\Psi:[0,1]\rightarrow\mathbb{R}$.
Take $\Psi$ to be the identity mapping, leading to direct regularized optimization of total preferences:
$\max\limits_{\pi_{\theta}}p^{*}(\pi\succ\mu)-\tau\mathbb{D}_{\textnormal{KL}}[\pi_{\theta}||\pi_{\textnormal{ref}}].$ (17)
To optimize the objective, IPO derives an off-policy expression:
$\min_{\pi_{\theta}}\underset{x,y_{w},y_{l},I\sim\mathcal{D}^{p}}{\mathbb{E}}\big{(}h_{\pi_{\theta}}(x,y_{w},y_{l},I)-\frac{\tau^{-1}}{2}\big{)}^{2},$ (18)
where $h_{\pi_{\theta}}$ is defined as:
$h_{\pi_{\theta}}(x,y_{w},y_{l},I)=I\cdot\log\big{(}\frac{\pi_{\theta}(y_{w}|x)\pi_{\textnormal{ref}}(y_{l}|x)}{\pi_{\theta}(y_{l}|x)\pi_{\textnormal{ref}}(y_{w}|x)}\big{)}+(1-I)\cdot\log\big{(}\frac{\pi_{\theta}(y_{l}|x)\pi_{\textnormal{ref}}(y_{w}|x)}{\pi_{\theta}(y_{w}|x)\pi_{\textnormal{ref}}(y_{l}|x)}\big{)}.$ (19)
IPO claims when preferences are deterministic or near deterministic, DPO will lead over-fitting to the preference dataset at the expense of ignoring the KL-regularation term. And IPO’s loss will always regularizes $\pi_{\theta}$ towards $\pi_{\textnormal{ref}}$ by controlling the gap between the log-likelihood ratios $\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}$ and $\log\frac{\pi_{\textnormal{ref}}(y_{w}|x)}{\pi_{\textnormal{ref}}(y_{l}|x)}$.
Similar to DPO, the IPO loss controls the ratio of $\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}$. So the optimal policy may be undefined, and also fails the KL-regularation term.
4 Method
In this work, we use the same objective of IPO. Unlike IPO’s loss, we derive a off-policy algorithm for preference maximization from a importance sampling policy optimization view, and add a intuitive sample loss to make KL-regularation term work.
4.1 Total Preference Optimization using Importance Sampling
First we define $x,y_{w},y_{l}$ as state, and $\mathcal{A}_{x,y_{w},y_{l}}=\{y_{w}\succ y_{l},y_{l}\succ y_{w}\}$ as action set, action as $\mathfrak{a}\in\mathcal{A}$ and define the reward of action $\mathfrak{a}$ as $r^{p}(\mathfrak{a})$, so
$\begin{split}r^{p}(\mathfrak{a}|x,y_{w},y_{l})&=r^{p}(\mathfrak{a}|x)\\ &=\mathbb{E}[\mathbb{I}\{\mathfrak{a}\}|x]\\ &=p^{*}(\mathfrak{a}|x).\end{split}$ (20)
Given a sample $x,y_{w},y_{l},I\in\mathcal{D}^{p}$, we can get rewards from both actions in $\mathcal{A}$:
$\{\langle x,y_{w},y_{l},I\rangle\}\rightarrow\{\langle\underbrace{(x,y_{w},y_{l})}_{\textnormal{state}},\underbrace{(y_{w}\succ y_{l})}_{\textnormal{action}},\underbrace{(I)}_{\textnormal{reward}}\rangle,\langle\underbrace{(x,y_{w},y_{l})}_{\textnormal{state}},\underbrace{(y_{l}\succ y_{w})}_{\textnormal{action}},\underbrace{(1-I)}_{\textnormal{reward}}\rangle\},$ (21)
and define the new dataset as $\bar{\mathcal{D}}^{p}$. Because both actions appear at the same time from $\bar{\mathcal{D}}^{p}$, we can define the policy generating $\bar{\mathcal{D}}^{p}$ as:
$\bar{\pi}^{p}(\mathfrak{a}|x,y_{w},y_{l})=\bar{\pi}^{p}(\mathfrak{a}|x)=1/2,\forall\mathfrak{a}\in\mathcal{A}.$ (22)
Now we can formulate the distribution to generate $\bar{\mathcal{D}}^{p}$ as:
$\begin{gathered}x\sim\rho\\ \langle y_{w},y_{l}\rangle\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\bar{\pi}^{p}(\cdot|x)\\ I\sim r^{p}(\mathfrak{a}|x)\end{gathered}.$ (23)
Define preference generating policy to be optimized as $\pi^{p}_{\theta}$, so the expected reward of $\pi^{p}_{\theta}$ over $\bar{\mathcal{D}}^{p}$ is:
$R(\pi^{p}_{\theta})=\underset{\begin{subarray}{c}x\sim\rho\\ \langle y_{w},y_{l}\rangle\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\pi^{p}_{\theta}(\cdot|x)\end{subarray}}{\mathbb{E}}[r^{p}(\mathfrak{a}|x)],$ (24)
and its easily to get $R(\bar{\pi}^{p})=1/2$.
Express the preferences maximization objective in reinforcement learning format:
$\max_{\pi^{p}_{\theta}}R(\pi^{p}_{\theta})$ (25)
According to REINFORCE algorithm, policy gradient of the 25 is:
$\underset{\begin{subarray}{c}x\sim\rho\\ \langle y_{w},y_{l}\rangle\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\pi^{p}_{\theta}(\cdot|x)\end{subarray}}{\mathbb{E}}[r^{p}(\mathfrak{a}|x)\log\pi^{p}_{\theta}(\mathfrak{a}|x)]$ (26)
Typically this gradient needs to be estimated from samples continually collected by $\pi^{p}_{\theta}$, which is data-inefficient. Using importance sampling, gradient 26 be expressed as:
$\underset{\begin{subarray}{c}x\sim\rho\\ \langle y_{w},y_{l}\rangle\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\bar{\pi}^{p}(\cdot|x)\end{subarray}}{\mathbb{E}}[\frac{\pi^{p}_{\theta}(\mathfrak{a}|x)}{\bar{\pi}^{p}(\mathfrak{a}|x)}r^{p}(\mathfrak{a}|x)\log\pi^{p}_{\theta}(\mathfrak{a}|x)]$ (27)
Gradient 27 can be calculated offline, and the corresponding algorithm is off-policy. According to equation 22, $\bar{\pi}^{p}(\mathfrak{a}|x)=1/2,\forall\mathfrak{a}\in\mathcal{A}$, so gradient 26 and 27 are identical.
Theorem 1
Objective 25 (which is identical to preferences maximization term of IPO) can be directly optimized using off-policy gradient ascent algorithm.
4.2 Off-policy Preference Learning under KL-regularation
Define the corresponding point-wise policy for $\pi^{p}_{\theta}$ as $\pi^{s}_{\theta}$, we use Bradley-Terry model to represent $\pi^{p}_{\theta}$ as:
$\pi^{p}_{\theta}(y_{w}\succ y_{l}|x)=\sigma(\log\pi^{s}_{\theta}(y_{w}|x)-\log\pi^{s}_{\theta}(y_{l}|x)).$ (28)
KL-regularized preferences maximization objective 25 can be expressed as:
$\begin{gathered}\max_{\pi^{p}_{\theta},\pi^{s}_{\theta}}R(\pi^{p}_{\theta})-\tau\mathbb{D}_{\textnormal{KL}}[\pi^{s}_{\theta}||\pi_{\textnormal{ref}}],\end{gathered}$ (29)
which includes two expectations over different distributions:
[…]
From theorem 1, preference maximization can be directly be solved with off-policy gradient descent. But the KL-regularization term needs to collect samples from $\pi^{s}_{\theta}(\cdot|x)$, which is on-policy. Additionally, to optimize KL-regularization with stochastic samples, it may ignore the regularization on $\pi_{\textnormal{ref}}(y_{w}|x),\pi_{\textnormal{ref}}(y_{l}|x)$.
Off-policy KL-regularation optimization
Minimize $\mathbb{D}_{\textnormal{KL}}[\pi^{s}_{\theta}||\pi_{\textnormal{ref}}]$ needs on-policy samples collection, which is data-inefficient. To solve the problem, we use use $\mathbb{D}_{\textnormal{KL}}[\pi_{\textnormal{ref}}||\pi^{s}_{\theta}]$ as the KL-regularization term. Because $\pi_{\textnormal{ref}}$ is fixed, we can generate samples $\mathcal{D}_{\pi_{\textnormal{ref}}(\cdot|x)}$ conditioned on $x\sim\rho$ in advance. Then when optimize $\mathbb{D}_{\textnormal{KL}}[\pi_{\textnormal{ref}}||\pi^{s}_{\theta}]$, we sample datas from $\mathcal{D}_{\pi_{\textnormal{ref}}(\cdot|x)}$.
Explicit regulariztion on $\pi_{\textnormal{ref}}(y_{w}|x),\pi_{\textnormal{ref}}(y_{l}|x)$
Optimize KL-regularization may using stochastic samples may ignore regularization on $\pi_{\theta}(y_{w}|x),\pi_{\theta}(y_{l}|x)$. To solve this problem, we add a KL-regularation term for preference probability.
Define $\pi^{p}_{\textnormal{ref}}(y_{w}\succ y_{l}|x)=\sigma(\log\pi_{\textnormal{ref}}(y_{w}|x)-\log\pi_{\textnormal{ref}}(y_{l}|x))$ , then
$\begin{split}\mathbb{D}_{\textnormal{KL}}[\pi^{p}_{\theta}||\pi^{p}_{\textnormal{ref}}]=&\underset{\begin{subarray}{c}x\sim\rho\\ y_{w},y_{l}\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\pi^{p}_{\theta}(\cdot|x)\end{subarray}}{\mathbb{E}}[\log\frac{\pi^{p}_{\theta}(\mathfrak{a}|x)}{\pi^{p}_{\textnormal{ref}}(\mathfrak{a}|x)}]\\ =&\underset{\begin{subarray}{c}x\sim\rho\\ y_{w},y_{l}\sim\mu(\cdot|x)\end{subarray}}{\mathbb{E}}[\sum_{\mathfrak{a}\in\mathcal{A}}\pi^{p}_{\theta}(\mathfrak{a}|x)\log\frac{\pi^{p}_{\theta}(\mathfrak{a}|x)}{\pi^{p}_{\textnormal{ref}}(\mathfrak{a}|x)}].\end{split}$ (31)
Maximum Preference Optimization (MPO) loss
Combine preference optimization term, point-wise policy KL-regularation term and pair-wise preference KL-regularation term, we get the final objective:
$\max_{\pi^{p}_{\theta},\pi^{s}_{\theta}}=\underset{x\sim\rho}{\mathbb{E}}\Big{[}\underset{y_{w},y_{l}\sim\mu(\cdot|x)}{\mathbb{E}}\big{\{}R(\pi^{p}_{\theta}(\cdot|x,y_{w},y_{l}))-\beta\mathbb{D}_{\textnormal{KL}}[\pi^{p}_{\theta}||\pi^{p}_{\textnormal{ref}}]\big{\}}-\tau\mathbb{D}_{\textnormal{KL}}[\pi_{\textnormal{ref}}||\pi^{s}_{\theta}]\Big{]}.$ (32)
With this objective, we define MPO loss on dataset $\bar{\mathcal{D}}^{p}$ and $\mathcal{D}_{\pi_{\textnormal{ref}}}$:
$\mathcal{L}_{\textnormal{MPO}}=\underset{\begin{subarray}{c}<x,y_{w},y_{l},\mathfrak{a},I>\sim\bar{\mathcal{D}}^{p}\\ \end{subarray}}{\mathbb{E}}\Big{[}-I\pi^{p}_{\theta}(\mathfrak{a}|x)+\beta\pi^{p}_{\theta}(\mathfrak{a}|x)\log\frac{\pi^{p}_{\theta}(\mathfrak{a}|x)}{\pi^{p}_{\textnormal{ref}}(\mathfrak{a}|x)}-\tau\underset{y\sim\mathcal{D}_{\textnormal{ref}}(\cdot|x)}{\mathbb{E}}[\log{\pi^{s}_{\theta}(y|x)}]\Big{]}.$ (33)
This loss is very simple and intuitive, term $-I\pi^{p}_{\theta}(\mathfrak{a}|x)$ try to maximize preferences, $\beta$ controls the strength of the KL regularization between pair-wise preference generating distribution, and $\tau$ controls the strength of the KL regularization between the point-wise generating distribution.
Eliminate $\pi_{\textnormal{ref}}$ after sample generation
When generating $\mathcal{D}_{\textnormal{ref}}(\cdot|x)$ and calculating $\log\pi^{p}_{\textnormal{ref}}(\mathfrak{a}|x)$ in advance, we can optimize $\mathcal{L}_{\textnormal{MPO}}$ without $\pi_{\textnormal{ref}}$. Which means we can optimize $\mathcal{L}_{\textnormal{MPO}}$ from $\pi_{\textnormal{ref}}$, without any modification of the model structure but loss function.