### 31 Dec 2023 v2

# Maximum Preference Optimization with Importance Sampling

Abstract
Preference learning is a key technology for aligning language models with human values. Reinforcement Learning from Human Feedback (RLHF) is a model based algorithm to optimize preference learning, which first fitting a reward model for preference score, and then optimizing generating policy with on-policy PPO algorithm to maximize the reward. The processing of RLHF is complex, time-consuming and unstable. Direct Preference Optimization (DPO) algorithm using off-policy algorithm to direct optimize generating policy and eliminating the need for reward model, which is data efficient and stable. DPO use Bradley-Terry model and log-loss which leads to over-fitting to the preference data at the expense of ignoring KL-regularization term when preference is deterministic. IPO uses a root-finding MSE loss to solve the ignoring KL-regularization problem. In this paper, we’ll figure out, although IPO fix the problem when preference is deterministic, but both DPO and IPO fails the KL-regularization term because the support of preference distribution not equal to reference distribution. Then, we design a simple and intuitive off-policy preference optimization algorithm from an importance sampling view, which we call Maximum Preference Optimization (MPO), and add off-policy KL-regularization terms which makes KL-regularization truly effective. The objective of MPO bears resemblance to RLHF’s objective, and likes IPO, MPO is off-policy. So, MPO attains the best of both worlds. To simplify the learning process and save memory usage, MPO eliminates the needs for both reward model and reference policy.
Keywords Large Language Model  $\cdot$ Preference Learning  $\cdot$ Reinforcement Learning
1 Introduction
Large language models (LLMs) Brown et al. (2020) Chowdhery et al. (2023) Bubeck et al. (2023) Radford et al. (2019) with massive scale parameters trained on large amount of data using pretrain, supervised fine-tune(SFT)Wei et al. (2021), and instruction fine-tune (IFT) Chung et al. (2022) algorithms has lead to surprising capabilities like in few-shot in context learning. The training data comes from variety of areas and has different quality, the training algorithms (pretrain, SFT, IFT) all based on maximum likelihood estimation (MLE) which learning to match the distribution of data. LLMs trained on these data using MLE algorithm generate contents with a quality gap to human judgement or values.
Preference learning Ziegler et al. (2019) Bai et al. (2022) Christiano et al. (2017) Stiennon et al. (2020) algorithm significantly improve generate quality to align with human values. It first collects pairs of generations under the same context, and a pairwise human preference to indicate which generation is better. Then a preference learning algorithm is used to optimize generating policy to generate better candidate from the pair.
Reinforcement learning form human feedback (RLHF)Ouyang et al. (2022) uses reward-model-based reinforcement learning algorithm to learn the optimal policy. It first learns a reward model from the preference data, then using an on-policy PPO Schulman et al. (2017) algorithm to maximize the learned reward. The reward is learned to use Bradley-Terry model Bradley and Terry (1952), which assumes the preference score can be approximated from substituted with point-wise reward. This assumption may lead to an approximation error when preference is deterministic. The PPO algorithm is used on data sampled from generating policy, which may has a different support or distribution drift from preference data, the learned reward model inference on the out-of-distribution data may reduce the accuracy. The process of RLHF need to train reward model and on-policy PPO algorithm which is complex, time-consuming, and unstable.
Direct preference optimization (DPO)Rafailov et al. (2023) combines off-policy algorithm and Bradley-Terry model to direct learn the generating policy from preference data. The off-policy algorithm based on KL-regularization reward maximization from off-RL community, which is data efficient, stable and eliminating the need for a reward model. When preference is deterministic which occurs in most cases, the reward of Bradley-Terry model is undefined, it leads to ignore the KL-regularization term and over-fitting the preference dataset.
Identity mapping preference optimization (IPO)Azar et al. (2023) also uses off-policy algorithm with KL-regularization to learn the generating policy from preference data. It learns to maximize preference probability under KL-regularization, and use root finding mean square error(MSE) loss to solve the maximization problem and fix the KL-regularization ignorance problem. But because of the support of preference data distribution is different from reference policy distribution, both DPO’s and IPO’s KL-regularization term fails.
In this paper, we design a simple and intuitive off-policy maximum preference optimization (MPO) algorithm from an importance sampling view, and add an off-policy KL-regularization term which makes KL-regularization truly effective. Our key contribution of this paper can summary as follows:
  • formalize preference learning as a preference/reward maximization problem, and design a simple and intuitive off-policy algorithm from importance sampling view
  • figure out KL-regularization fails when optimized on preference data, and design an off-policy sample loss to make KL-regularation truly effective
  • eliminate the reward substitution assumption and out-of-distribution generalization assumption
  • eliminate the needs for both reward model and reference policy to save memory usage
Figure 1: Maximum Preference Optimization (MPO) direct optimize preference maximization on preference data using off-policy algorithm, and use offline SFT, pretrain data to make KL-regularation truly effective, which also eliminate the needs for both reward model and reference policy.
2 Preliminaries
The main pipeline of preference learning usually consists of three phases: 1) pretraining and supervised fine-tuning (SFT), where SFT is not a must; 2) preference data collection; 3) reinforcement-learning optimization.
Pretraining and SFT phase
Preference learning typically started with a pretrained LLMs or LLMs fine-tuned on high quality data using maximum likelihood estimation. We define the final policy after this phase as $\pi_{\text{ref}}$, and the data to train $\pi_{\text{ref}}$ as $\mathcal{D}_{\text{ref}}$, so,
$\pi_{\text{ref}}\approx\arg\max_{\pi}\mathbb{E}_{x,y\sim\mathcal{D}_{\text{ref}}}\log\pi(x)\log(y|x)$ (1)
Preference data collection phase
After pretraining and SFT phase, $\pi_{\text{ref}}$ is prompted by context $x$, and generate two responses $y_{w},y_{l}\sim\pi_{\textnormal{ref}}(\cdot|x)$. Then $x,y_{w},y_{l}$ is labeled by humans to judge which response is preferred, and denote $y_{w}\succ y_{l}|x$ if $y_{w}$ is preferred, and $y_{l}\succ y_{w}|x$ if $y_{l}$ is preferred. We define a new symbol $I=\mathbb{I}[y_{w}\succ y_{l}|x]$, and all $<x,y_{w},y_{l},I>$ consist the preference dataset $\mathcal{D}^{p}$:
$\langle x,y_{w},y_{l},I\rangle\sim\mathcal{D}^{p}.$ (2)
We also define $\rho$ as the context distribution of $x$ and $\mu$ as the preference pair distribution given context $x$ from preference data distribution.
$x\sim\rho$ (3)
$y_{w},y_{l},I\sim\mu.$ (4)
Let $p^{*}(y_{w}\succ y_{l}|x)=\mathbb{E}_{y_{w},y_{l},I\sim\mu}[{\mathbb{I}\{I=1\}}|x]$, which denotes the preference probability of $y_{w}\succ y_{l}$ given context $x$. Then the expected preference of $y_{w}$ over $\mu$, noted $p^{*}(y\succ\mu|x)$, via the following equation:
$p^{*}(y\succ\mu|x)=\underset{y_{l}\sim\mu(\cdot|x)}{\mathbb{E}}[p^{*}(y_{w}\succ y_{l}|x)].$ (5)
For any policy $\pi$, denote the total preference of policy $\pi$ to $\mu$ as
$p^{*}(\pi\succ\mu)=\underset{x\sim\rho\atop y_{\sim}\pi(\cdot|x)}{\mathbb{E}}[p^{*}(y\succ\mu|x)].$ (6)
Reinforcement-learning optimization phase
At the final phase, the prevailing method is using reinforcement learning algorithm to learn an explicit or implicit reward from the preference data, and then using on-policy or off-policy policy gradient algorithm to maximize the reward. Recently, some methods derive the optimal policy using reward maximization under KL-regularization and also derive a loss with optimal policy as its solution, then learn the optimal policy by minimizing the derived loss on empirical dataset.
3 Background
3.1 Reinforcement Learning from Human Feedback (RLHF)
In this paper, we define a new pair-wise preference reward $r^{p}(y_{w}\succ y_{l}|x)=p^{*}(y_{w}\succ y_{l}|x)$, and design new algorithm to direct optimize the preference(reward) maximum objective.
The RLHF uses standard two-phase reward-model based reinforcement learning to maximize the reward. It contains two step: 1) reward estimation from preference data 2) reward maximization using PPO algorithm.
Reward estimation from preference data
In previous research, the point-wise reward is learned to use Bradley-Terry model. Given context x, define $r^{*}(y|x)$ as the reward of generating $y$. Bradley-Terry model assumes the preference probability $p^{*}(y_{w}\succ y_{l}|x)$ as:
$p^{*}(y_{w}\succ y_{l}|x)=\frac{\exp(r^{*}(y_{w}|x))}{\exp(r^{*}(y_{w}|x))+\exp(r^{*}(y_{l}|x))}=\sigma(r^{*}(y_{w}|x)-r^{*}(y_{l}|x))$ (7)
where $\sigma(\cdot)$ is the sigmoid function.
RLHF uses 7 to model the point-wise reward, and optimize log loss to estimate the reward. The estimated reward is defined as parameterized $r_{\phi}$, and loss function defined as:
$\mathcal{L}(r_{\phi})=-\underset{x,y_{w},y_{l},I\sim\mathcal{D}^{p}}{\mathbb{E}}[I\cdot\sigma(r_{\phi}(y_{w}|x)-r_{\phi}(y_{l}|x))+(1-I)\cdot\sigma(r_{\phi}(y_{l}|x)-r_{\phi}(y_{w}|x))],$ (8)
where $I=\mathbb{I}[y_{w}\succ y_{l}|x]$.
The loss 8 is doing maximize likelihood estimation, and the estimated reward $r_{\phi}$ is used to approximate probability $p(y_{w}\succ y_{l}|x)$ from preference data distribution $\mathcal{D}^{p}$.
Reward maximization using PPO algorithm
The reward-maximization or KL-regularized reward-maximization objective is used for reinforcement learning policy optimization:
$\max_{\pi_{\theta}}\underset{{y\sim\pi_{\theta}(\cdot|x)}}{\mathbb{E}}[r_{\phi}(y|x)]-\beta\mathbb{D}_{\text{KL}}[\pi_{\theta}||\pi_{\text{ref}}]$ (9)
where $\mathbb{D}_{\text{KL}}$ is the KL-divergence and $\beta$ is the regularization weight. This objective is optimized by on-policy REINFORCE Mnih et al. (2016) or PPO algorithm.
The second phase of RLHF is optimizing objective 9 using $r_{\phi}$ learned from 8. PPO is an on-policy algorithm which continues collect data from the current policy $\pi_{\theta}$ and estimated reward model, then it uses these data to estimate the gradient of 9, and then update the current policy. Because $\pi_{\theta}$ is different from $\pi_{\text{ref}}$ defined as 1, samples generated by $\pi_{\theta}$ has a different distribution of $D^{p}$. So, RLHF assumes $r_{\phi}$ can generalize to out-of-distribution samples generated by $\pi_{\theta}$. Following prior work Rafailov et al. (2023), it is straightforward to show that the optimal solution of 9 takes the form:
$\pi_{\theta}\propto\pi_{\text{ref}}\exp(\frac{1}{\beta}r_{\phi}(x,y)).$ (10)
RLHF also mixing a pretraining gradient into the PPO objective, in order to fix the performance regression on public NLP datasets. And RLHF call the final objective ’PPO-ptx’. Define $\mathcal{D}_{\text{pretrain}}$ as the pretraining dataset, then the combined objective defined as:
$\max_{\pi_{\theta}}\underset{\langle x,y\rangle\sim\pi_{\theta}}{\mathbb{E}}\bigl{[}r_{\theta}(x,y)]-\beta\log(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)})\bigr{]}+\gamma\underset{\mathcal{D}_{\text{pretrain}}}{\mathbb{E}}[\pi_{\theta}(x)].$ (11)
3.2 Direct Preference Optimization (DPO)
An alternative RL algorithm to preference learning is direct preference optimization (DPO), which eliminates the training of reward model. DPO derives a reward from Eq. 10:
$r(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}+\beta\log Z(x),$ (12)
where $Z(x)=\sum_{y}\pi_{\text{ref}}(y|x)\exp(\frac{1}{\beta}r(x,y))$ is the partition function. Substituting the re-parameterized reward in 12 int to the Bradley-Terry model 7:
$\begin{split}p_{\theta}(y_{w}\succ y_{l}|x)&=\sigma(\beta h_{\pi_{\theta}}(x,y_{w},y_{l}))\\ &=\frac{1}{1+\exp\big{(}\beta h_{\pi_{\theta}}(x,y_{w},y_{l})\big{)}}\end{split},$ (13)
where $h_{\pi_{\theta}}(x,y_{w},y_{l})$ is defined as:
$\log\frac{\pi_{\theta}(y_{w}|x)\pi_{\text{ref}}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)\pi_{\theta}(y_{l}|x)}.$ (14)
Substituting the probability 13 to the log loss 8, DPO formulates a maximum likelihood objective for the parameterized policy $\pi_{\theta}$ given the empirical preference dataset $\mathcal{D}^{p}$:
$\begin{split}\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\underset{x,y_{w},y_{l},I\sim\mathcal{D}^{p}}{\mathbb{E}}\bigl{[}I&\sigma(\beta h_{\pi_{\theta}}(x,y_{w},y_{l}))+(1-I)\sigma(\beta h_{\pi_{\theta}}(x,y_{w},y_{l}))\bigr{]}.\end{split}$ (15)
Although this pair-wise loss eliminates the need to calculate the partition $Z(x)$, it also makes the optimal solution $\pi^{*}_{\theta}$ undefined when there are not enough constraints. For example, if weight $\pi_{\theta}(y_{w}|x)$ and $\pi_{\theta}(y_{l}|x)$ with the same multiplier $M$, the logits of the sigmoid function $\sigma$ will remain the same:
$\beta\log\frac{\pi_{\theta}(y_{w}|x)*M}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)*M}{\pi_{\text{ref}}(y_{l}|x)}=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}.$ (16)
This will make the final learned policy $\pi_{\theta}$ suboptimal, and also fails the KL-regularation term.
3.3 $\Psi$-PO with dientity mapping (IPO)
IPO defines a new objective called $\Psi$-preference optimization objective ($\Psi$PO):
$\max\limits_{\pi_{\theta}}\underset{\begin{subarray}{c}x\sim\rho\\ y_{w}\sim\pi_{\theta}(\cdot|x)\\ y_{l}\sim\mu(\cdot|x)\end{subarray}}{\mathbb{E}}[\Psi(p^{*}(y_{w}\succ y_{l}|x))]-\tau\mathbb{D}_{\text{KL}}[\pi_{\theta}||\pi_{\text{ref}}],$ (17)
where $\Psi$ is a general non-decreasing function $\Psi:[0,1]\rightarrow\mathbb{R}$.
Take $\Psi$ to be the identity mapping, leading to direct regularized optimization of total preferences:
$\max\limits_{\pi_{\theta}}p^{*}(\pi\succ\mu)-\tau\mathbb{D}_{\text{KL}}[\pi_{\theta}||\pi_{\text{ref}}].$ (18)
To optimize the objective, IPO derives an off-policy loss on empirical dataset:
$\min_{\pi_{\theta}}\underset{x,y_{w},y_{l},I\sim\mathcal{D}^{p}}{\mathbb{E}}\bigl{[}Ih_{\pi_{\theta}}(x,y_{w},y_{l})-(1-I)h_{\pi_{\theta}}(x,y_{l},y_{w})-\frac{\tau^{-1}}{2}\bigr{]}^{2},$ (19)
IPO claims when preferences are deterministic or near deterministic, DPO will lead over-fitting to the preference dataset at the expense of ignoring the KL-regularation term. And IPO’s loss will always regularizes $\pi_{\theta}$ towards $\pi_{\text{ref}}$ by controlling the gap between the log-likelihood ratios $\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}$ and $\log\frac{\pi_{\text{ref}}(y_{w}|x)}{\pi_{\text{ref}}(y_{l}|x)}$.
Similar to DPO, the IPO loss controls the ratio of $\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}$, not $\pi_{\theta}(y_{w}|x)$ nor $\pi_{\theta}(y_{l}|x)$. When there are not enough constraints, which in all most all cases, the optimal policy is undefined, so also fails the KL-regularation term.
4 Method
In this work, we combine the preference maximization term of IPO’s loss and modified regularization term of RLHF’s loss. Unlike IPO, which derive the off-policy from preference maximization under KL-regularization, we formulate preference maximization as a reward maximization problem in reinforcement learning setting, and derive an off-policy objective from an importance sampling based policy optimization view, but without the help of KL-regularation. Then we combine the off-policy reward maximization objective with modified regularization terms of RLHF’s ’PPO-ptx’ objective, which makes the KL-regularization truly effective. We call the algorithm Maximum Preference Optimization (MPO). The final objective of MPO bears resemblance to RLHF’s objective, and likes IPO, MPO is off-policy.
4.1 Preference(reward) Maximization with Importance Sampling
We define preference as reward, and formalize preference maximization as reward maximization problem in reinforcement learning setting. Define $x,y_{w},y_{l}$ as state, and $\mathcal{A}_{x,y_{w},y_{l}}=\{y_{w}\succ y_{l},y_{l}\succ y_{w}\}$ as action set, action as $\mathfrak{a}\in\mathcal{A}$ and define the reward of action $\mathfrak{a}$ as $r^{p}(\mathfrak{a}|x,y_{w},y_{l})$, which is the preference probability. We simplify $r^{p}(\mathfrak{a}|x,y_{w},y_{l})$ as $r^{p}(\mathfrak{a}|x)$, so:
$\begin{split}r^{p}(\mathfrak{a}|x,y_{w},y_{l})&=r^{p}(\mathfrak{a}|x)\\ &=\mathbb{E}[\mathbb{I}\{\mathfrak{a}\}|x]\\ &=p^{*}(\mathfrak{a}|x).\end{split}$ (20)
Given a sample $x,y_{w},y_{l},I\in\mathcal{D}^{p}$, we can get rewards from both actions in $\mathcal{A}$:
$\{\langle x,y_{w},y_{l},I\rangle\}\rightarrow\{\langle\underbrace{(x,y_{w},y_{l})}_{\text{state}},\underbrace{(y_{w}\succ y_{l})}_{\text{action}},\underbrace{(I)}_{\text{reward}}\rangle,\langle\underbrace{(x,y_{w},y_{l})}_{\text{state}},\underbrace{(y_{l}\succ y_{w})}_{\text{action}},\underbrace{(1-I)}_{\text{reward}}\rangle\},$ (21)
and define the converted dataset as $\bar{\mathcal{D}}^{p}$. Because both actions appear at the same time from $\bar{\mathcal{D}}^{p}$, we can define the policy generating $\bar{\mathcal{D}}^{p}$ as:
$\bar{\pi}^{p}(\mathfrak{a}|x,y_{w},y_{l})=1/2,\forall\mathfrak{a}\in\mathcal{A}.$ (22)
Given the state $\langle x,y_{w},y_{l}\rangle$, we simplify $\bar{\pi}^{p}(\mathfrak{a}|x,y_{w},y_{l})$ as $\bar{\pi}^{p}(\mathfrak{a}|x)$. Now we can formulate the distribution to generate $\bar{\mathcal{D}}^{p}$ as:
$\begin{gathered}x\sim\rho\\ \langle y_{w},y_{l}\rangle\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\bar{\pi}^{p}(\cdot|x)\\ I\sim r^{p}(\mathfrak{a}|x)\end{gathered}.$ (23)
Define preference generating policy to be optimized as $\pi^{p}_{\theta}$, so the expected reward of $\pi^{p}_{\theta}$ over $\bar{\mathcal{D}}^{p}$ is:
$R(\pi^{p}_{\theta})=\underset{\begin{subarray}{c}x\sim\rho\\ \langle y_{w},y_{l}\rangle\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\pi^{p}_{\theta}(\cdot|x)\end{subarray}}{\mathbb{E}}[r^{p}(\mathfrak{a}|x)],$ (24)
and it’s easy to get $R(\bar{\pi}^{p})=1/2$.
Express the preference maximization objective in reinforcement learning context:
$\max_{\pi^{p}_{\theta}}R(\pi^{p}_{\theta}).$ (25)
Typically, the gradient of the objective 25 needs to be estimated from samples continually collected by $\pi^{p}_{\theta}$, which is data-inefficient. But, for preference maximization, we can directly estimate gradient from dataset $\bar{\mathcal{D}}^{p}$, which is off-policy.
Theorem 1.
Preference(Reward) maximization objective 25 (which is identical to preferences maximization term of IPO) can be directly optimized using off-policy gradient ascent algorithm.
Proof.
According to REINFORCE algorithm, policy gradient of the 25 is:
$\nabla R(\pi^{p}_{\theta})=\underset{\begin{subarray}{c}x\sim\rho\\ \langle y_{w},y_{l}\rangle\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\pi^{p}_{\theta}(\cdot|x)\end{subarray}}{\mathbb{E}}[r^{p}(\mathfrak{a}|x)\log\pi^{p}_{\theta}(\mathfrak{a}|x)].$ (26)
Using importance sampling, gradient 26 be expressed as:
$\nabla R(\pi^{p}_{\theta})=\underset{\begin{subarray}{c}x\sim\rho\\ \langle y_{w},y_{l}\rangle\sim\mu(\cdot|x)\\ \mathfrak{a}\sim\bar{\pi}^{p}(\cdot|x)\end{subarray}}{\mathbb{E}}[\frac{\pi^{p}_{\theta}(\mathfrak{a}|x)}{\bar{\pi}^{p}(\mathfrak{a}|x)}r^{p}(\mathfrak{a}|x)\log\pi^{p}_{\theta}(\mathfrak{a}|x)]$ (27)
Gradient 27 can be calculated offline, and the corresponding algorithm is off-policy. According to equation 22, $\bar{\pi}^{p}(\mathfrak{a}|x)=1/2,\forall\mathfrak{a}\in\mathcal{A}$, so gradient 26 and 27 are identical. ∎
4.2 Off-policy Preference Learning under KL-regularation
Define the corresponding point-wise policy for $\pi^{p}_{\theta}$ as $\pi^{s}_{\theta}$, we use Bradley-Terry model to approximate $\pi^{p}_{\theta}$ as:
$\pi^{p}_{\theta}(y_{w}\succ y_{l}|x)=\sigma(\log\pi^{s}_{\theta}(y_{w}|x)-\log\pi^{s}_{\theta}(y_{l}|x)).$ (28)
Then, reward maximization objective 25 can be expressed:
$\max_{\pi^{s}_{\theta}}R(\pi^{p}_{\theta}),$ (29)
which means we optimize $\pi^{s}_{\theta}$ to maximize the preference reward for corresponding policy $\pi^{p}_{\theta}$.
Let $\pi^{s}_{\text{ref}}=\pi_{\text{ref}}$, like RLHF’s ’PPO-ptx’ objective 11, KL-regularized preferences maximization objective 25 can be expressed as:
$\begin{gathered}\max_{\pi^{s}_{\theta}}R(\pi^{p}_{\theta})-\beta\mathbb{D}_{\text{KL}}[\pi^{s}_{\theta}||\pi^{s}_{\text{ref}}]+\gamma\underset{x\sim\mathcal{D}_{\text{pretrain}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(x)]\end{gathered}.$ (30)
From theorem 1, preference maximization term $R(\pi^{p}_{\theta})$ can be directly be solved with off-policy policy gradient method. Pretraining data regularization term $\mathbb{E}_{s\sim\mathcal{D}_{\text{pretrain}}}\log\pi^{s}_{\theta}(x)$ can also be solved with offline data. But the KL-regularization term $\mathbb{D}_{\text{KL}}[\pi^{s}_{\theta}||\pi^{s}_{\text{ref}}]$ needs to collect samples from $\pi^{s}_{\theta}(\cdot|x)$, which is on-policy.
Off-policy KL-regularation on reference policy $\pi^{s}_{\text{ref}}$
Minimize $\mathbb{D}_{\text{KL}}[\pi^{s}_{\theta}||\pi^{s}_{\text{ref}}]$ needs on-policy samples collection, which is data-inefficient. To solve the problem, we replace $\mathbb{D}_{\text{KL}}[\pi^{s}_{\theta}||\pi^{s}_{\text{ref}}]$ with $-\mathbb{E}_{\langle x,y\rangle\sim\mathcal{D}_{\text{ref}}}[\log\pi^{s}_{\theta}(y|x)]$. Like pretraining data regularization, replaced regularization can be computed with offline data.
Maximum Preference Optimization (MPO) loss
With the modified regularization on $\pi^{s}_{\text{ref}}$, we get the final objective of MPO:
$\max_{\pi^{s}_{\theta}}R(\pi^{p}_{\theta})-\beta\underset{\langle x,y\rangle\sim\mathcal{D}_{\text{ref}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(y|x)]-\gamma\underset{x\sim\mathcal{D}_{\text{pretrain}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(x)]$ (31)
With this objective, we define empirical MPO loss on dataset $\bar{\mathcal{D}}^{p}$,$\mathcal{D}_{\text{ref}}$ and $\mathcal{D}_{\text{pretrain}}$:
$\mathcal{L}_{\text{MPO}}=\underset{\begin{subarray}{c}<x,y_{w},y_{l},\mathfrak{a},I>\sim\bar{\mathcal{D}}^{p}\\ \end{subarray}}{\mathbb{E}}[-I\pi^{p}_{\theta}(\mathfrak{a}|x)]+\beta\underset{\langle x,y\rangle\sim\mathcal{D}_{\text{ref}}}{\mathbb{E}}[\log\pi^{s}_{\theta}(y|x)]+\gamma\underset{y\sim\mathcal{D}_{\text{pretrain}}}{\mathbb{E}}[\log{\pi^{s}_{\theta}(y|x)}].$ (32)
This loss is very simple and intuitive, term $-I\pi^{p}_{\theta}(\mathfrak{a}|x)$ try to maximize preferences, $\beta$ controls the strength of the regularization on SFT dataset, and $\gamma$ controls the strength of the regularization pretraining dataset.
Eliminate both the need for reward model and $\pi_{\text{ref}}$
By using preference as reward, we don’t need a reward model to approximate preference probability. By replacing KL-regularization $\mathbb{D}_{\text{KL}}[\pi^{s}_{\theta}||\pi^{s}_{\text{ref}}]$ with offline dataset regularization $-\mathbb{E}_{\langle x,y\rangle\sim\mathcal{D}_{\text{ref}}}[\log\pi^{s}_{\theta}(y|x)]$, we don’t need the reference policy $\pi^{s}_{\text{ref}}$. So, MPO algorithm simplify the learning process and saves memory usage.